### v6

@@ 1 Introduction
Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (1.1a)
@@
where the initial value $v$ is given and $W_{H}$ is an infinite-dimensional Wiener process.
@@
  • To the best of our knowledge, the existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may encompass randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis when compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [17, 18]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates employing the more general spatial $L^{q}$-norms are relatively scarce.
Accordingly, the main contributions of this work are as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are pivotal for the numerical analysis. Specifically, we affirm that
@@
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [28, Theoerem 3.2].
@@
To achieve a second-order spatial convergence rate, the previous numerical analysis in the literature typically enforce stringent regularity assumptions on the diffusion term. However, in this study, we impose only some moderate regularity conditions on $F$ (see Section 2). Furthermore, this work introduces a novel approach to the numerical analysis of finite element method-based spatial semidiscretizations for a broad class of nonlinear stochastic parabolic equations.
The rest of this paper is organized as follows. Section 2 introduces some conventions and definitions used in this paper, and the spatial semidiscretization analyzed in this research. Section 3 presents the main results of this paper, including the regularity of Eq. 1.1 with non-smooth initial data and the convergence of the spatial semidiscretization under both non-smooth and smooth initial conditions. Section 4 provides a rigorous proof of the aforementioned results. Lastly, Section 5 summarizes the key findings of this work.
The rest of this paper is organized as follows. Section 2 establishes the foundational notation, definitions, and preliminaries that underpin the analysis within this study, as well as detailing the spatial semidiscretization that is the focus of our investigation. Section 3 is dedicated to the principal findings of this work, which encompass the treatment of the regularity of the solution to (1.1) with nonsmooth initial data, along with the convergence analysis of the spatial semidiscretization scheme under both nonsmooth and smooth initial conditions. Section 4 furnishes a detailed and exacting proof of the results previously outlined. Finally, Section 5 offers a concise synthesis of the pivotal insights and contributions that this research brings to the field.
@@ 2 Preliminaries
Notational Conventions. Throughout this paper, we will adhere to the following notational conventions. For Banach spaces $E_{1}$ and $E_{2}$, we denote the space of all bounded linear operators from $E_{1}$ to $E_{2}$ as $\mathcal{L}(E_{1},E_{2})$, and it is abbreviated to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity map is denoted by $I$. For any $\theta\in(0,1)$ and $p\in(1,\infty)$, we use $(\cdot,\cdot)_{\theta,p}$ to represent the real interpolation space constructed by the $K$-method, as described in [21, Chapter 1]. The conjugate exponent of $p$ is denoted by $p^{\prime}$. For $q\in[1,\infty]$, we briefly write $L^{q}(\mathcal{O})$ as $L^{q}$. Let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces (see, e.g., [14]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. For $p\in(1,\infty)$ and a Banach space $X$, we use $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$ to denote the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ denotes the space of all $\gamma$-radonifying operators from $H$ to $X$ (see [11, Chapter 9]). Finally, for each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H,$
where $H$ is a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
Definition of Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, satisfying the following conditions for every $n\in\mathbb{N}$:
@@
Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying
@@
Consequently, the model equation Eq. 1.1 can be explicitly stated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
Spatial Semidiscretization. Let $\mathcal{K}_{h}$ represent a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplexes. We define the finite element space $X_{h}$ as
$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\,\middle|\,v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\text{ and }v_{h}\text{ vanishes on }\partial\mathcal{O}\right\},$
where $C(\overline{\mathcal{O}})$ is the space of continuous functions over the closure of $\mathcal{O}$, $r\geqslant 1$ is an integer denoting the polynomial degree, and $P_{r}(K)$ is the space of polynomials defined on simplex $K$ up to degree $r$. Let $h$ denote the maximum diameter of the elements in $\mathcal{K}_{h}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and we define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ by
@@
where $\mu$ represents the three-dimensional Lebesgue measure on $\mathcal{O}$. The spatial semidiscretization considered in this work is given by
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,$ (2.7a)
@@
where $v$ is the given initial value.
Mild Solutions. Let $p\in(1,\infty)$ and $q\in[2,\infty)$. We define the convolution operator $S_{0}$ acting on any function $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T.$ (2.8)
Similarly, for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, we introduce the stochastic convolution operator $S_{1}$ defined by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T.$ (2.9)
An $\mathbb{F}$-adapted process $y$ is called a mild solution to Eq. 1.1 if $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ and satisfies
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.10)
where $G$ is defined as
$G(t):=e^{t\Delta}v,\quad t\in[0,T].$ (2.11)
In the context of the spatial semidiscretization, for any $g_{h}\in L^{p}(0,T;X_{h})$, we define
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}t,\quad t\in[0,T].$
For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,X_{h})))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
An $\mathbb{F}$-adapted $X_{h}$-valued stochastic process $y_{h}$ is called a mild solution to the spatial semidiscretization Eq. 2.7 if $y_{h}\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);X_{h})$ and it satisfies that
$y_{h}=G_{h}+S_{0,h}(y_{h}-P_{h}y_{h}^{3})+S_{1,h}P_{h}F(y_{h}),$
where $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (2.12)
@@ 3 Main results
In the remainder of this paper, the symbol $c$ represents a generic positive constant that depends solely on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{h}$, the regularity parameters of the triangulation $\mathcal{K}_{h}$, the indices of the pertinent Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ in Eqs. 1.1 and 2.7, and the exponents related to the spatial mesh size $h$. The value of $c$ may differ from line to line. In addition, for any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
@@
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
@@
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
@@
For Theorem 3.1, we have the following remarks.
@@
    Liu [17] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields the result that
    provided that $v\in\dot{H}^{1,2}$. However, the theory in [17] does not cover the case of non-smooth initial data. In the context of the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [4, Lemma 3.1] provide some a priori estimates with the general spatial $L^{q}$ norms.
@@
    Assume that the condition Eq. 2.1 is removed. By an approximation argument we conclude that Eq. 3.1 remains valid. Moreover, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, the following assertions hold:
    $y-G\in L^{p}(\Omega\times(0,T);\dot{H}^{1,q}),$
    $y-G\in L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p})).$
Remark 3.2.
For Theorems 3.2 and 3.3, we have the following remarks.
  1. 1.
    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative infinite-dimensional Wiener process, the existing literature typically imposes stringent regularity conditions on the diffusion coefficient. Our chosen formulation of the function $F$ adheres to the criteria specified in Assumption 2.2 of Liu and Qiao’s work [20]; however, it fails to satisfy the more stringent requirements imposed by their Assumption 2.3. Examining the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, particularly the relaxation of condition (2.1), holds significant theoretical and practical implications.
  2. 2.
    In Section 4.3, we demonstrate that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, provided that $v$ is an element of $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.51). It can be rigorously shown that this result remains valid even if $v$ is an element of $W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This conclusion is crucial for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, the question of whether this conclusion extends to the case where $v$ is an element of $L^{\infty}$ remains unresolved.

### v7

@@ 1 Introduction
Let $\mathcal{O}$ be a bounded convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator with homogeneous Dirichlet boundary conditions. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
@@
where the initial value $v$ is given, and $W_{H}$ is an infinite-dimensional $H$-cylindrical Brownian motion.
@@
  • To the best of our knowledge, existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may include randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [18, 19]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates using the more general spatial $L^{q}$-norms are relatively scarce.
The main contributions of this work are as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are crucial for the numerical analysis. Specifically, we affirm that
@@
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [35, Theoerem 3.2].
@@
To achieve a second-order spatial convergence rate, prior numerical analyses in the literature have typically enforced stringent regularity assumptions on the diffusion coefficients. In contrast, this study imposes relatively moderate regularity conditions on the diffusion coefficient $F$ (see Section 2). While our numerical analysis assumes deterministic initial data, extending the results to accommodate random initial data is straightforward.
The remainder of this paper is organized as follows. Section 2 introduces the notational conventions and definitions that are used throughout the paper. Section 3 presents the main results of this work, including the regularity properties of the mild solution to model problem 1.1 with non-smooth initial data, and establishes the convergence of the finite element-based spatial semidiscretization for both non-smooth and smooth initial conditions. A rigorous proof of these results is provided in Section 4. Finally, Section 5 concludes the paper by summarizing the key findings of this study.
@@ 2 Preliminaries
Notational Conventions. Throughout this paper, we will use the following conventions. For Banach spaces $E_{1}$ and $E_{2}$, the space of all bounded linear operators from $E_{1}$ to $E_{2}$ is denoted by $\mathcal{L}(E_{1},E_{2})$, and we abbreviate this to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity operator is indicated by $I$. Given $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ is used to refer to the real interpolation space constructed via the $K$-method, as detailed in Chapter 1 of [22]. The conjugate exponent of each $p\in[0,\infty]$ is denoted by $p^{\prime}$. The symbol $\mathcal{O}$ represents a bounded convex domain in $\mathbb{R}^{3}$ with a boundary $\partial\mathcal{O}$ that belongs to the $C^{2,0}$ class (see Section 6.2 of [11]). For $q\in[1,\infty]$, we abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$, and use $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ to denote two standard Sobolev spaces (see, e.g., [34]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is defined by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
For $p\in(1,\infty)$ and a Banach space $X$, the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$ is denoted by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ represents the space of all $\gamma$-radonifying operators from $H$ to $X$, as defined in [13, Chapter 9]. For a comprehensive treatment of the theory of stochastic integration with respect to the $H$-cylindrical Wiener process $W_{H}$, we refer the reader to [24].
Definition of the Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, which satisfy the following conditions for every $n\in\mathbb{N}$:
@@
Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the summability condition
@@
Accordingly, the model equation (1.1) can be explicitly formulated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0\leqslant t\leqslant T,\\ y(0)=v.\end{cases}$
Mild Solutions. Let $p,q\in[2,\infty)$. We introduce the deterministic convolution operator $S_{0}$ acting on functions $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad\text{for }0\leqslant t\leqslant T.$ (2.7)
Additionally, for any process $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, we define the stochastic convolution operator $S_{1}$ by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad\text{for }0\leqslant t\leqslant T.$ (2.8)
Here, $e^{t\Delta}$, for $t\in[0,\infty)$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ subject to homogeneous Dirichlet boundary conditions. A process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ is called a mild solution to equation (1.1) if it satisfies the equality
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.9)
where the function $G$ is given by
$G(t):=e^{t\Delta}v,\quad\text{for }t\in[0,T].$ (2.10)
Note that the conditions in the above definition of a mild solution can be relaxed (see, e.g., [29, Chapter 6]). However, the present formulation is adequate for the purposes of this study.
Spatial Semidiscretization. Consider $\mathcal{K}_{h}$ to be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, composed of three-dimensional simplexes. We introduce the finite element space $X_{h}$ as
$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\bigm{|}v_{h}=0\text{ on }(\mathcal{O}\setminus\mathcal{O}_{h})\cup\partial\mathcal{O}_{h},\text{ and }v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Bigr{\}},$
where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is a positive integer, and $P_{r}(K)$ is the space of polynomials on simplex $K$ up to degree $r$. Let the spatial mesh size $h$ be the maximum diameter of the elements in $\mathcal{K}_{h}$. Define $P_{h}$ as the $L^{2}$-orthogonal projection operator onto $X_{h}$. The discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ is defined by
@@
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. We consider the following spatial semidiscretization:
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (2.11a)
@@
where $v$ is the given initial value. Assume that $v\in L^{2}$. Since this spatial semidiscretization is a finite-dimensional stochastic differential equation (SDE), we can use the classical theory of finite-dimensional SDEs (see, e.g., [28, Theorem 3.21]) to conclude that Eq. 2.11 has a unique strong solution.
@@ 3 Main results
In the remainder of this paper, the symbol $c$ denotes a generic positive constant that depends only on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{n}$, the constant $C_{F}$ in Eqs. 2.1, 2.2, 2.4 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ given in equations Eq. 1.1 and Eq. 2.11, the regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with the spatial mesh size $h$. The value of $c$ may differ at each occurrence. For any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\|v\|_{\dot{H}^{\alpha,q}}:=\|(-\Delta)^{\alpha/2}v\|_{L^{q}}\quad\text{for all $v\in\dot{H}^{\alpha,q}$}.$
The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$. It is well-established that $\Delta$ has a unique extension as a bounded linear operator from $\dot{H}^{\alpha,q}$ to $\dot{H}^{\alpha-2,q}$ for all $\alpha\in[0,2)$. Specifically, when $\alpha=2$, the space $\dot{H}^{2,q}$ is identical to the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$ and is endowed with a norm equivalent to that of $W^{2,q}(\mathcal{O})$ (see [11, Theorem 9.15]). Moreover, for $\alpha=1$, the space $\dot{H}^{1,q}$ coincides with the standard Sobolev space $W_{0}^{1,q}(\mathcal{O})$, and the two norms are equivalent (see [38, Theorem 16.14]).
@@
Assume that $v\in L^{\infty}$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq. 2.11. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
@@
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution of Eq. 2.11. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
@@
It is worth noting that the aforementioned theoretical results remain valid under the assumption that the operator $F$ possesses the growth and Lipschitz continuity properties outlined in Lemma 4.5.
@@
Liu [18] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields
provided that $v\in\dot{H}^{1,2}$. Furthermore, Liu and Qiao [21] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem. Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
Remark 3.2.
For Theorems 3.2 and 3.3, we have the following comments.
@@
    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, the existing literature often imposes strict regularity conditions on the diffusion coefficients. For instance, the study in [23] demands that the diffusion coefficients possess bounded second-order derivatives. Our choice of diffusion coefficient $F$ adheres to the criteria outlined in Assumption 2.2 of [21]; however, it does not meet the more stringent requirements stated in Assumption 2.3 of the same reference. Investigating the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, especially the loosening of condition (2.1), has important theoretical and practical implications.
@@
    In Section 4.3, we show that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, assuming $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.53). It can be rigorously demonstrated that this result also holds when $v\in W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This finding is critical for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, it remains an open question whether this conclusion applies when $v\in L^{\infty}$.